# A Spectral Localization Method for Time-Fractional Integro-Differential Equations with Nonsmooth Data

### v1

@@ 3.2 Error Analysis in Temporal Semi-discrete CIM Scheme
The estimate of $\|g(\phi,t)\|_{L^{2}(\Omega)}$ was given in Theorem 1. For the truncation error $TE$, we have
$TE$ $=\tau\sum_{k=N}^{+\infty}\|g(k\tau,t)\|_{L^{2}(\Omega)}$
$\leq\tau\sum_{k=N}^{\infty}\Bigl{(}C\varphi(\theta,\tilde{c})\,e^{\mu t}\bigl{(}\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\bigr{)}e^{-\mu t\sin(\theta-\tilde{c})\cosh(k\tau)}\Bigr{)}$
$\leq C\varphi(\theta,\tilde{c})\,e^{\mu t}\bigl{(}\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\bigr{)}\cdot\int_{N\tau}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh x}\,dx.$
By Lemma 3,
$\int_{N\tau}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh x}\,dx\leq\Bigl{(}1+L\bigl{(}\mu t\sin(\theta-\tilde{c})\bigr{)}\Bigr{)}e^{-\mu t\sin(\theta-\tilde{c})\cosh(N\tau)},$
@@
$TE\leq C\varphi(\theta,\tilde{c})L\bigl{(}\mu t\sin(\theta-\tilde{c})\bigr{)}\bigl{(}\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\bigr{)}e^{\mu t}\cdot\left(\frac{1}{e^{\mu t\sin(\theta-\tilde{c})\cosh(N\tau)}}\right).$ (3.7)
As for the disretization error $DE$, according to the estimate (3.2) and Lemma 3, it follows that
$DE$ $\leq\frac{\int_{-\infty}^{+\infty}\|g(\nu+i\tilde{c},t)\|_{L^{2}(\Omega)}\,d\nu}{e^{2\pi\tilde{c}/\tau}-1}$ (3.8)
$\leq C\varphi(\theta,\tilde{c})\,e^{\mu t}\bigl{(}\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\bigr{)}\bigl{(}\int_{-\infty}^{+\infty}|e^{-\mu t\sin(\theta-\tilde{c})\cosh x}|\,dx\bigr{)}\cdot\left(\frac{1}{e^{2\pi\tilde{c}/\tau}-1}\right)$
$=C\varphi(\theta,\tilde{c})\,e^{\mu t}\bigl{(}\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\bigr{)}\bigl{(}2\int_{0}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh x}\,dx\bigr{)}\cdot\left(\frac{1}{e^{2\pi\tilde{c}/\tau}-1}\right)$
$\leq C\varphi(\theta,\tilde{c})\,L\bigr{(}\mu t\sin(\theta-\tilde{c})\bigl{)}\bigl{(}\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\bigr{)}e^{\mu t}\cdot\left(\frac{1}{e^{2\pi\tilde{c}/\tau}-1}\right).$
@@
Eq. (3.9) reveals that for $t\in[t_{0},t_{1}]$, $t_{1}=\Lambda t_{0}$, and $\Lambda\geq 1$, the truncation error $TE$ decreases as time $t$ grows. Therefore, our strategy for selecting parameters is to equilibrate the truncation error at the initial time $t=t_{0}$ and the discretization error $DE$. Specifically, we enforce the condition
@@
while simultaneously minimizing $DE$. Solving Eq. (3.10) yeilds
@@
Solving the system given by Eq. (3.11) and (3.12) yields the explicit expressions
$\tau=\frac{1}{N}\rm{arcosh}\left(\frac{\Lambda}{(1-\eta)\sin(\theta-\tilde{c})}\right):=\frac{P(\eta)}{N},\quad\mu=\frac{2\pi\tilde{c}(1-\eta)N}{\Lambda t_{0}P(\eta)}.$
@@
$\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}$
$\leq$ $DE+TE$
$\leq$ $C\varphi(\theta,\tilde{c})L\bigl{(}\mu t_{0}\sin(\theta-\tilde{c})\bigr{)}\left(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\right)\times e^{\mu\Lambda t_{0}}\left(\frac{1}{e^{2\pi\tilde{c}/\tau}-1}+\frac{1}{e^{\mu t_{0}\sin(\theta-\tilde{c})\cosh(N\tau)}}\right)$
$\leq$ $C\varphi(\theta,\tilde{c})L\bigl{(}\mu t_{0}\sin(\theta-\tilde{c})\bigr{)}\left(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\right)\times\left(\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\right).$
@@
In fact, as demonstrated in numerical results, with an appropriate parameter choice, the actual error of the CIM exhibits the asymptotic behavior
@@
    Spectral Accuracy: Noting that $P(\eta)>\rm{arcosh}^{-1}(\Lambda)=\ln(\Lambda+\sqrt{\Lambda^{2}-1})\sim\ln(\Lambda)$, we confirm the method’s spectral accuracy:
    $\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}=O\left(e^{-CN}\right)\quad\text{with}\quad C=O(1/\ln\Lambda).$ (3.13)
@@
    Long-Time Simulation: Eq. (3.13) reveals a critical trade-off: a larger $\Lambda$ (i.e., a longer simulation time $t_{1}=\Lambda t_{0}$) results in a smaller constant $C$, leading to a slower exponential decay. Consequently, achieving a prescribed accuracy for long-time simulations requires a significantly larger number of quadrature points $N$. This inherent characteristic of the CIM motivates the development of the acceleration algorithms discussed previously in Subsection 2.4.
@@
by maximizing the function $Q(\eta):=2\pi\tilde{c}\eta/P(\eta)$. Denote $Q(\eta*):=\max_{0<\eta<1}{Q(\eta)}$. Substituting $\eta^{*}$ into Eq. (3.6) then yields the optimal parameters $\mu^{*}$ and $\tau^{*}$ for the contour $\Gamma$. The complete parameter selection procedure is presented as follows:
Input: $\theta$; $t_{0}$; $\Lambda\geq 1$; $t_{1}=\Lambda t_{0}$; NN.
@@
1. $P(\eta)\leftarrow\rm{arcosh}\left(\frac{\Lambda}{(1-\eta)\sin(\theta-\bar{c})}\right)$;
@@ 3. $\eta^{*}\leftarrow\max_{\eta\in(0,1)}Q(\eta)$;
4. $\tau^{*},\mu^{*}\leftarrow$ Substitute $\eta^{*}$ into Eq. (3.6) to get the optimal parameters $\tau^{*}$ and $\mu^{*}$.
@@ 3.3 Error Analysis of the Fully Discrete Scheme
The following two lemmas provide error estimates for the spatial semi-discrete scheme (2.14), addressing the nonhomogeneous problem with zero initial data and the homogeneous problem with non-smooth initial data, respectively. The proofs can be found in Theorems 3.1 and 3.2 of [22].
@@
Assume $f=0$ and $u_{0}\in L^{2}(\Omega)$. Let $u(t)$ and $u_{h}(t)$ be the solutions given by (2.2) and (2.14), respectively. Then for $u_{0,h}=P_{h}u_{0}$, we have
@@
Assume $f\neq 0$ and $u_{0}=0$. Let $u(t)$ and $u_{h}(t)$ be the solutions given by (2.2) and (2.14), respectively. Then
$\|u(t)-u_{h}(t)\|_{L^{2}(\Omega)}+h\|u(t)-u_{h}(t)\|_{H_{0}^{1}(\Omega)}\leq Ch^{2}\left(\|f(0)\|+\int_{0}^{t}\|f^{\prime}(s)\|\,ds\right).$
@@
Let $u(t)$ and $u_{h}^{N}(t)$ be the solutions given by (2.2) and (2.15), respectively. Consider the countour (2.8), with $z\in\Sigma_{\tilde{\vartheta}}$, where $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ for $0<\varepsilon<1$, and let the parameters $\mu$ and $\tau$ be determined according to (3.6). For $t\in[t_{0},t_{1}]$, with $t_{1}=\Lambda t_{0}$ and $\Lambda\geq 1$, if $f=0$ and $u_{0}\in L^{2}(\Omega)$, then the following error estimate holds:
$\|u(t)-u_{h}^{N}(t)\|_{L^{2}(\Omega)}\leq C\left(h^{2}t^{-\beta}+\varphi(\theta,\tilde{c})L\bigl{(}\mu t_{0}\sin(\theta-\tilde{c})\bigr{)}\left(\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\right)\right)\|u_{0}\|_{L^{2}(\Omega)}.$
@@
Let $u(t)$ and $u_{h}^{N}(t)$ be the solutions given by (2.2) and (2.15), respectively. Consider the countour (2.8), with $z\in\Sigma_{\tilde{\vartheta}}$, where $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ for $0<\varepsilon<1$, and let the parameters $\mu$ and $\tau$ be determined according to (3.6). For $t\in[t_{0},t_{1}]$, with $t_{1}=\Lambda t_{0}$ and $\Lambda\geq 1$, if $f\neq 0$ and $u_{0}=0$, then the following error estimate holds:
@@
$\leq$ $C\left(h^{2}\left(\|f(0)\|+\int_{0}^{t}\|f^{\prime}(s)\|ds\right)+\varphi(\theta,\tilde{c})L\bigl{(}\mu t_{0}\sin(\theta-\tilde{c})\bigr{)}\left(\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\right)\|\widehat{f}(z)\|_{S}\right).$
@@ 4 Numerical Results
This section presents we present four numerical examples to demonstrate the efficacy of the proposed spectral localization method for time-fractional integro-differential equations. The domain is set to $\Omega=(0,1)$ or $(0,1)\times(0,1)$, and the final time is $T=1$. For temporal discretization, we employ the CIM discussed in this paper. Specifically, for the hyperbolic contour (2.8), we select the parameter $\theta=0.6767$ and set the initial time as $t_{0}=0.1$. Spatial discretization is carried out using the standard Galerkin FEM.
Given a fixed spatial mesh size $h$, the temporal errors are defined as
$\text{Err}_{\tau}(N):=\max_{t_{0}\leq t\leq\Lambda t_{0}}\left\|u(\cdot,t)-u_{h}^{N}(\cdot,t)\right\|_{L^{2}(\Omega)},$
@@
$\text{Err}_{h}(t):=\left\|u(\cdot,t)-u_{h}^{N}(\cdot,t)\right\|_{L^{2}(\Omega)}.$
@@
$\text{Order}:=\frac{\ln{(\mathrm{Err}_{h}/Err_{\frac{h}{2}})}}{\ln 2}.$
All numerical examples were computed in MATLAB R2020a on a system with an Intel(R) Core(TM) i5-1035G1 CPU @ 1.00GHz.

### v2

@@ 3.2 Error Analysis in Temporal Semi-discrete CIM Scheme
Applying the estimate of $\|g(\phi,t)\|_{L^{2}(\Omega)}$ given in Theorem 1, we have the estimate of the truncation error $TE=\|I_{\tau}-I_{\tau;N}\|_{L^{2}(\Omega)}$ as follows
$TE$ $\leq\tau\sum_{k=N}^{+\infty}\|g(\phi_{k},t)\|_{L^{2}(\Omega)}$
$\leq\tau\sum_{k=N}^{\infty}\big(C\varphi(\theta,\tilde{c})\,e^{\mu t}(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S})e^{-\mu t\sin(\theta-\tilde{c})\cosh((k+\tfrac{1}{2})\tau)}\big)$
$\leq C\varphi(\theta,\tilde{c})\,e^{\mu t}\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)\cdot\int_{N\tau}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh x}\,dx.$
By Lemma 3, it follows
$\int_{N\tau}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh x}\,dx\leq\big(1+L(\mu t\sin(\theta-\tilde{c}))\big)e^{-\mu t\sin(\theta-\tilde{c})\cosh(N\tau)},$
@@
$TE\leq C\varphi(\theta,\tilde{c})L\big(\mu t\sin(\theta-\tilde{c})\big)\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)e^{\mu t}\cdot\frac{1}{e^{\mu t\sin(\theta-\tilde{c})\cosh(N\tau)}}.$ (3.7)
As for the discretization error $DE=\|I-I_{\tau}\|_{L^{2}(\Omega)}$, it follows from the estimate (2.6) and Lemma 3 that
$DE$ $\leq\frac{1}{e^{2\pi\tilde{c}/\tau}-1}\int_{-\infty}^{+\infty}\|g(\nu+i\tilde{c},t)\|_{L^{2}(\Omega)}\,d\nu$ (3.8)
$\leq C\varphi(\theta,\tilde{c})\,e^{\mu t}\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)\Big(\int_{-\infty}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu}\,d\nu\Big)\cdot\frac{1}{e^{2\pi\tilde{c}/\tau}-1}$
$=C\varphi(\theta,\tilde{c})\,e^{\mu t}\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)\Big(2\int_{0}^{+\infty}e^{-\mu t\sin(\theta-\tilde{c})\cosh\nu}\,d\nu\Big)\cdot\frac{1}{e^{2\pi\tilde{c}/\tau}-1}$
$\leq C\varphi(\theta,\tilde{c})\,L\big(\mu t\sin(\theta-\tilde{c})\big)\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)e^{\mu t}\cdot\frac{1}{e^{2\pi\tilde{c}/\tau}-1}.$
@@
It indicates in (3.9) that the truncation error $TE$ decreases as time $t$ grows for $t\in[t_{0},t_{1}]$, $t_{1}=\Lambda t_{0}$, and $\Lambda\geq 1$. Therefore, our strategy for selecting parameters is to equilibrate the truncation error at the initial time $t=t_{0}$ and the discretization error $DE$. Specifically, we enforce the condition
@@
while simultaneously minimizing $DE$. Solving (3.10) yields
@@
Solving the system given by (3.11) and (3.12) yields the explicit expressions
$\tau=\frac{1}{N}{\rm{arcosh}}\Big(\frac{\Lambda}{(1-\eta)\sin(\theta-\tilde{c})}\Big):=\frac{P(\eta)}{N},\quad\mu=\frac{2\pi\tilde{c}(1-\eta)N}{\Lambda t_{0}P(\eta)}.$
@@
$\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}\leq DE+TE$
$\leq C\varphi(\theta,\tilde{c})L\big(\mu t_{0}\sin(\theta-\tilde{c})\big)\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)e^{\mu\Lambda t_{0}}\Big(\frac{1}{e^{2\pi\tilde{c}/\tau}-1}+\frac{1}{e^{\mu t_{0}\sin(\theta-\tilde{c})\cosh(N\tau)}}\Big)$
$\leq C\varphi(\theta,\tilde{c})L\big(\mu t_{0}\sin(\theta-\tilde{c})\big)\big(\|u_{0}\|_{L^{2}(\Omega)}+\|\widehat{f}(z)\|_{S}\big)\Big(\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\Big).$
@@
In fact, as demonstrated in numerical results, with an appropriate parameter choice, the actual error of the temporal semi-discrete CIM scheme (2.12) exhibits the asymptotic behavior
@@
    Spectral Accuracy: Noting that $P(\eta)>\rm{arcosh}(\Lambda)=\ln(\Lambda+\sqrt{\Lambda^{2}-1})\sim\ln(\Lambda)$, we confirm the method’s spectral accuracy:
    $\|u(t)-u^{N}(t)\|_{L^{2}(\Omega)}=O(e^{-CN})\quad\text{with}\quad C=O(1/\ln\Lambda).$ (3.13)
@@
    Long-Time Simulation: (3.13) reveals a critical trade-off: a larger $\Lambda$ (i.e., a longer simulation time $t_{1}=\Lambda t_{0}$) results in a smaller constant $C$, leading to a slower exponential decay. Consequently, achieving a prescribed accuracy for long-time simulations requires a significantly larger number of quadrature points $N$. This inherent characteristic of the CIM motivates the development of the acceleration algorithms discussed previously in Subsection 2.4.
@@
by maximizing the function $Q(\eta):=2\pi\tilde{c}\eta/P(\eta)$. Denote $Q(\eta*):=\max_{0<\eta<1}{Q(\eta)}$. Substituting $\eta^{*}$ into (3.6) then yields the optimal parameters $\mu^{*}$ and $\tau^{*}$ for the contour $\Gamma$. The complete parameter selection procedure is presented as follows.
Input: $\theta$, $t_{0}$, $\Lambda\geq 1$, $N$.
@@
1. $P(\eta)\leftarrow\rm{arcosh}\Big(\frac{\Lambda}{(1-\eta)\sin(\theta-\bar{c})}\Big)$;
@@ 3. $\eta^{*}\leftarrow\max_{\eta\in(0,1)}Q(\eta)$;
4. $\tau^{*},\mu^{*}\leftarrow$ Substitute $\eta^{*}$ into (3.6) to get the optimal parameters $\tau^{*}$ and $\mu^{*}$.
@@ 3.3 Error Analysis of the Fully Discrete Scheme
The following two lemmas provide error estimates for the spatial semi-discrete scheme (2.16), addressing the nonhomogeneous problem with zero initial data and the homogeneous problem with nonsmooth initial data, respectively. The proofs can be found in Theorems 3.1 and 3.2 of [22].
@@
Assume $f\equiv 0$ and $u_{0}\in L^{2}(\Omega)$. Let $u(t)$ and $u_{h}(t)$ be the solutions given by (2.2) and (2.16), respectively. Then for $u_{0,h}=P_{h}u_{0}$, we have
@@
Assume $f\neq 0$ and $u_{0}\equiv 0$. Let $u(t)$ and $u_{h}(t)$ be the solutions given by (2.2) and (2.16), respectively. Then
$\|u(t)-u_{h}(t)\|_{L^{2}(\Omega)}+h\|u(t)-u_{h}(t)\|_{H_{0}^{1}(\Omega)}\leq Ch^{2}\Big(\|f(0)\|+\int_{0}^{t}\|f^{\prime}(s)\|ds\Big).$
@@
Let $u(t)$ and $u_{h}^{N}(t)$ be the solutions given by (2.2) and (2.17), respectively. Consider the contour (2.10), with $z\in\Sigma_{\tilde{\vartheta}}$, where $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ for $0<\varepsilon<1$, and let the parameters $\mu$ and $\tau$ be determined according to (3.6). For $t\in[t_{0},t_{1}]$, with $t_{1}=\Lambda t_{0}$ and $\Lambda\geq 1$, if $f\equiv 0$ and $u_{0}\in L^{2}(\Omega)$, then the following error estimate holds:
$\|u(t)-u_{h}^{N}(t)\|_{L^{2}(\Omega)}\leq C\Big(h^{2}t^{-\beta}+\varphi(\theta,\tilde{c})L\big(\mu t_{0}\sin(\theta-\tilde{c})\big)\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\Big)\|u_{0}\|_{L^{2}(\Omega)}.$
@@
Let $u(t)$ and $u_{h}^{N}(t)$ be the solutions given by (2.2) and (2.17), respectively. Consider the contour (2.10), with $z\in\Sigma_{\tilde{\vartheta}}$, where $\alpha+\beta\in(\varepsilon,2-\varepsilon)$ for $0<\varepsilon<1$, and let the parameters $\mu$ and $\tau$ be determined according to (3.6). For $t\in[t_{0},t_{1}]$, with $t_{1}=\Lambda t_{0}$ and $\Lambda\geq 1$, if $f\neq 0$ and $u_{0}\equiv 0$, then the following error estimate holds:
@@
$\leq C\Big(h^{2}\Big(\|f(0)\|+\int_{0}^{t}\|f^{\prime}(s)\|ds\Big)+\varphi(\theta,\tilde{c})L\big(\mu t_{0}\sin(\theta-\tilde{c})\big)\frac{e^{\mu\Lambda t_{0}}}{e^{2\pi\tilde{c}/\tau}-1}\|\widehat{f}(z)\|_{S}\Big).$
@@ 4 Numerical Results
In this section, we present four numerical examples to demonstrate the efficacy of the proposed spectral localization method for time-fractional integro-differential equations. The domain is set to $\Omega=(0,1)$ or $(0,1)\times(0,1)$. For temporal discretization, we employ the CIM discussed in this paper. Specifically, for the hyperbolic contour (2.10), we select the parameter $\theta=0.6767$ and set the initial time as $t_{0}=0.1$. $T=\Lambda t_{0}$ for some given $\Lambda$. All the other parameters are determined in accordance with Algorithm 1. Spatial discretization is carried out using the standard piecewise linear Galerkin FEM.
Given a fixed spatial mesh size $h$, the temporal errors are defined by
$\text{Err}_{\tau}(N):=\max_{t_{0}\leq t\leq\Lambda t_{0}}\|u(\cdot,t)-u_{h}^{N}(\cdot,t)\|_{L^{2}(\Omega)},$
@@
$\text{Err}_{h}(t):=\|u(\cdot,t)-u_{h}^{N}(\cdot,t)\|_{L^{2}(\Omega)}.$
@@
$\text{Order}:=\frac{\ln{\big(\mathrm{Err}_{h}(t)/\mathrm{Err}_{\frac{h}{2}}(t)\big)}}{\ln 2}.$
All numerical examples were computed in MATLAB R2020a on a laptop with an Intel(R) Core(TM) i5-1035G1 CPU @ 1.00GHz.